OpenStudy (acespeedfighter):

What is the value of 1/2 x^3 + 3.4y when x = 4 and y = 2? (4 points) 12.8 14.8 25.8 38.8

OpenStudy (_citrus_):

\[\frac{ 1 }{ 2 }4^3 +3.4(2)\] Multiply 3.4 and 2 = 6.8 4 cubed= 64 because 4*4*4 And then reduce the fraction by dividing numerator and denominator by 2= 1*32 Gives you 38.8

OpenStudy (acespeedfighter):

Oh ok i see

OpenStudy (acespeedfighter):

but wait if

OpenStudy (acespeedfighter):

oh never mind i see it

OpenStudy (_citrus_):

:)

OpenStudy (acespeedfighter):

hx